The profit function for the Haggard Apple Orchard`s out-of-state sales is
P(x) = -x squared + 276x - 11,454, where x is the number of pounds of apples in hundreds and P(x) is the profit in dollars. How many pounds should be sold out of state to produce a maximum profit, and what is that profit?




Answer by stanbon(57953)

P(x) = -x squared + 276x - 11,454, where x is the number of pounds of apples in hundreds and P(x) is the profit in dollars. How many pounds should be sold out of state to produce a maximum profit, and what is that profit?
-------------------
P(x) = -x squared + 276x - 11,454
P(x) = -x^2 + 276x - 11454
---
max occurs when x = -b/2a = -276/(2*-1) = 138
---------------
Sell 138 lbs. out-of-state to maximize profit.
===========================
Cheers,
Stan H.